OpenStudy (anonymous):

12x2 + 9x2 = do you add the exponents or do they stay the same

OpenStudy (anonymous):

no for 2 reasons

OpenStudy (anonymous):

so its 21x2

OpenStudy (anonymous):

1 never add exponents in addition. 2 you can add exponents if the bases are the same. not only are the bases not the same (2) but it is not even multiplication

OpenStudy (anonymous):

yup :D

hero (hero):

You don't add exponents because x^2 factors out: 12x^2 + 9x^2 = (12 + 9)x^2 = 21x^2

OpenStudy (anonymous):

9x + 8x = like is this one would it be just 17x

hero (hero):

exactly